base of temple is a rectangle with a perimeter of 300m. the area of the base is 4400m. what are the dimensions of the base?
all i have is
l+w=perimeter
lxw=4400

Answer by checkley77(12844)

L*W=4400
L=4400/W
2L+2W=300
2(4400/W)+2W=300
(8800+2W*W)/W=300
8800+2W^2=300W
2W62-300W+8800=0
2(W^2-150W+4400)
2(W-110)(W-40)=0
W-110=0
W=110 ANSWER FOR THE WIDTH & LENGTH=40
OR LENGTH=40 & WIDTH=10
PROOF:
40*110=4400
4400=4400
2*110+2*40=300
220+80=300
300=300